I have a word problem that I cannot seem to figure out.
One number is 12 more than another. The sum of the smaller number and twice the larger is 39. Find the larger number.
No matter how I try to work out this problem it comes out wrong. Please Help!!! Answer by ankor@dixie-net.com(22740) 
 One number is 12 more than another. The sum of the smaller number and twice the larger is 39. Find the larger number.
:
Let x = "one number the (larger)"; let y = "another" number
:
Write a simple expression for each sentence:
:
"One number is 12 more than another"
x = 12 + y
:
"sum of the smaller number and twice the larger is 39"
y + 2x = 39
:
1st equation states that x = (12 + y), so substitute [y+12] for x in the 2nd eq
y + 2(y+12) = 39
y + 2y + 24 = 39
3y = 39 - 24
y = 15/3
y = 5
:
x = 12 + 5
x = 17
:
Check:
"sum of the smaller number and twice the larger is 39."
5 + 2(17) = 39